Vase I contains 2 white and 4 black balls and Vase II contains 4 white and 4 black balls. If a ball is drawn out at random from one of the two vases, what is the probability that it is a white ball?

Grade:12th pass

2 Answers

Vikas TU
6 years ago
Dear Student,
probability of selecting one vase=1/2
          Probability of selecting white ball from vase 1 = 2/6
          Probability of selecting white ball from vase 2 = 4/8
         Therefore,  Probability of selecting white ball at randim from one of the two vases= (1/2)x(2/6) + (1/2)x(4/8)
                          = 1/6 + ¼
                          = 5/12
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Dheeraj Upadhyay
4 years ago
The probability of first one will be

 

no.of white balls / total no. of balls

 

i.e. 2/6=1/3

 

the probability of second one will be:

 

no. of white balls / total no. of balls

 

i.e. 4/8

 

=1/2

 

total probability :

 

no.of white balls / total no. of balls

 

i.e. 6/14

 

=3/7